what is the solution for y=x-2 2x+2y=4
Answer by mananth(16949)

x-y= 2 .............1
2x+2y=4
/2
x+y=2 .............2

x-y= 2
x+y= 2

Add the two equations
2x=4

/2

x=2

plug value of x in (1)

x-y= 2
2-y= 2
y=0